# CodeCrash: Stress Testing LLM Reasoning under Structural and Semantic Perturbations

### v1

@@ B.3 Inserting Garbage Code
Listing 8: Garbage death loop function templates for GBC perturbation.
@@
Listing 9: GBC-perturbed example from LCB (sample 36) for output prediction.
@@
B.4 Aggregating Structural Perturbations
As introduced in §3.3.1, we aggregate the application of REN, RTF, and GBC to a single code snippet, as shown in Code 10.
Listing 10: ALL-perturbed example from LCB (sample 36) for output prediction.
@@ B.4 Aggregating Structural Perturbations
B.5 Inserting Misleading Messages
As introduced in §3.3.2, we insert misleading comments for key AST node types, including function declarations, return statements, loops, conditionals, variable assignments, and common data structure operations. The data structure operations include common list, set, and string manipulations: append, extend, insert, remove, pop, sort, reverse, update, add, split, join, replace, lower, upper, capitalize, and swap case. To construct these comments, we design prompts for GPT-4o to generate misleading but plausible messages for each type. In this experiment, we manually selected 20 messages for each AST node type and 5 for each data structure operation. Take variable assignments as an example; Code 11 is the prompt for generating misleading comments for variable assignments, and Code 12 is a list of selected misleading comment candidates. Code 13 and 14 present an example of MDC and MPS perturbations.
Listing 11: Prompt used to generate misleading comments for variable assignments.
@@ B.5 Inserting Misleading Messages
Listing 12: Selected misleading comments for variable assignments used in MDC and MPS perturbations, where ‘{variable}’ is the variable name.
@@
Listing 13: MDC-perturbed example from LCB (sample 36) for output prediction.
@@
Listing 14: MPS-perturbed example from LCB (sample 36) for output prediction.
@@
B.6 Inserting Incorrect Hints
In order to generate plausible, logical but incorrect hints, we instruct GPT-4o to rewrite the correct assertion expression to incorrect ones and only allow it to modify the output value using the prompt in Code 15 for output prediction. An example perturbed instance is shown in Code 16, where the ground-truth return value 2 is replaced with the misleading hint # The return value is 3.
In contrast, we use the prompt shown in Code 17 to generate misleading input hints, allowing GPT-4o to perform CoT reasoning before generating the expression. Since some functions return the same output regardless of the input arguments, we instruct GPT-4o to force the expression to become incorrect by modifying input arguments, such as changing data types, swapping parameter order, or adding/removing arguments. An example perturbed instance is shown in Code 18.
Listing 15: Prompt for generating misleading hint comments for output prediction.
@@ B.6 Inserting Incorrect Hints
Listing 16: Prompt for generating misleading hint comments for input prediction.
@@
Listing 17: Prompt for generating misleading hint comments for input prediction.
@@
Listing 18: MHC-perturbed example from CRUX (sample 2) for input prediction.
@@
Appendix C Supplementary Experimental Results
C.1 Impact on Output Prediction
@@
Table 8: Relative PASS@1 degradation of LLMs under all perturbation types for output prediction tasks across CRUX and LCB in direct inference (N=3) [TO ABSOLUTE SCORE].
@@
C.2 Robustness Across Model Scales
In §4.1.3, we compare the robustness against different model scales. To fairly aggregate model robustness across different perturbation types, we compute a weighted average of the relative PASS@1 degradation under four representative perturbations: ALL (structural), MDC and MPS (misleading messages), and MHC (hints). We select ALL as the representative structural perturbation because of the compounding effects of REN, RTF, and GBC captured in §4.1.1, which reflect the most challenging structural setting. Since MDC and MPS stem from the same perturbation strategy and often target overlapping AST node types, we assign them lower weights ($1/8$ each) to avoid overrepresenting similar effects. Overall, we define a weighted robustness metric with weights of $1/2$, $1/8$, $1/8$, and $1/4$ for ALL, MDC, MPS, and MHC, respectively, and summarize the results in Table 2. For completeness, we also report the results under uniform weighting ($1/4$ for each) in Table 9, which show similar trends and confirm the stability of our findings in §4.1.3.
Table 9: Perturbations Output Direct (Relative)
@@ C.2 Robustness Across Model Scales
C.3 Effect of CoT Reasoning
Table 10: Comparison of absolute PASS@1 degradation aggregated over CRUX and LCB for output prediction tasks in direct inference (N=3) and CoT inference (N=1) [TO RELATIVE SCORE].

### v2

@@ A.3 Inserting Garbage Code
Code 4: Garbage death loop function templates for GBC perturbation.
@@
Code 5: GBC-perturbed example from LCB (sample 36) for output prediction.
@@
A.4 Aggregating Structural Perturbations
As introduced in §3.3.1, we aggregate the application of REN, RTF, and GBC to a single code snippet, as shown in Code 6.
Code 6: ALL-perturbed example from LCB (sample 36) for output prediction.
@@ A.4 Aggregating Structural Perturbations
A.5 Inserting Misleading Messages
As introduced in §3.3.2, we insert misleading comments for key AST node types, including function declarations, return statements, loops, conditionals, variable assignments, and common data structure operations. The data structure operations include common list, set, and string manipulations: append, extend, insert, remove, pop, sort, reverse, update, add, split, join, replace, lower, upper, capitalize, and swap case. To construct these comments, we design prompts for GPT-4o to generate misleading but plausible messages for each type. In this experiment, we manually selected 20 messages for each AST node type and 5 for each data structure operation. Take variable assignments as an example; Code 7 is the prompt for generating misleading comments for variable assignments, and Code 8 is a list of selected misleading comment candidates. Code 9 and 10 present an example of MDC and MPS perturbations.
Code 7: Prompt used to generate misleading comments for variable assignments.
@@ A.5 Inserting Misleading Messages
Code 8: Selected misleading comments for variable assignments used in MDC and MPS perturbations, where ‘{variable}’ is the variable name.
@@
Code 9: MDC-perturbed example from LCB (sample 36) for output prediction.
@@
Code 10: MPS-perturbed example from LCB (sample 36) for output prediction.
@@
A.6 Inserting Incorrect Hints
In order to generate plausible, logical but incorrect hints, we instruct GPT-4o to rewrite the correct assertion expression to incorrect ones and only allow it to modify the output value using the prompt in Code 11 for output prediction. An example perturbed instance is shown in Code 12, where the ground-truth return value 2 is replaced with the misleading hint # The return value is 3.
In contrast, we use the prompt shown in Code 13 to generate misleading input hints, allowing GPT-4o to perform CoT reasoning before generating the expression. Since some functions return the same output regardless of the input arguments, we instruct GPT-4o to force the expression to become incorrect by modifying input arguments, such as changing data types, swapping parameter order, or adding/removing arguments. An example perturbed instance is shown in Code 14.
Code 11: Prompt for generating misleading hint comments for output prediction.
@@ A.6 Inserting Incorrect Hints
Code 12: Prompt for generating misleading hint comments for input prediction.
@@
Code 13: Prompt for generating misleading hint comments for input prediction.
@@
Code 14: MHC-perturbed example from Crux (sample 2) for input prediction.
@@
Appendix B Statistical Analysis
Due to resource constraints, we generate three candidates for direct inference per query. To assess the reliability of this setup, we compute the standard deviation ($\sigma$) across each perturbation and vanilla setting.
We present two summary tables for output prediction tasks (Table 8) and input prediction tasks (Table 9). The results show that across a total of $357$ measurements, only 9 cases have $\sigma>1.0$ (i.e., $0.01$ on a scale from 00 to $1$), and $57$ have $\sigma>0.5$, indicating that most scores exhibit low variance under repeated sampling.
Furthermore, our benchmark includes $1279$ problems across $8$ experiment settings on $17$ models for output prediction tasks, and $800$ problems across $5$ experiment settings on $17$ models for input prediction tasks. Among $241,944$ combinations, $231,726$ have identical correctness labels across all three candidates, performing an overall stability rate of $\mathbf{95.78\%}$.
@@ Appendix B Statistical Analysis
Table 8: Standard deviation of LLM performance under all perturbation types for output prediction tasks across Crux and LCB in direct inference mode ($N=3$). Dark red highlights denote $\sigma>1.0$, and light red highlights denote $\sigma>0.5$.
Model Series Model Name VAN REN RTF GBC ALL MDC MPS MHC
Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB
GPT-4 Omni GPT-4o-Mini $0.2125$ $0.1705$ $0.2125$ $0.429$ $0.27$ $0.7873$ $0.5303$ $0.1968$ $0.2569$ $0.429$ $0.27$ $0.1705$ $0.1768$ $0.0984$ $0.2569$ $0.5905$
GPT-4o $0.4125$ $0.6818$ $0.27$ $0.5114$ $0.3584$ $0.2952$ $0.2041$ $0.858$ $0.4082$ $0.4921$ $0.6374$ $0.451$ $0.598$ $0.451$ $0.2125$ $0.8056$
Claude Series Claude-3.5-Haiku-20241022 $1.546$ $0.8523$ $1.562$ $6.987$ $1.373$ $1.45$ $0.6208$ $0.5986$ $0.2041$ $0.1705$ $0.2041$ $0.6146$ $0.3864$ $0.2604$ $0.1021$ $0.2604$
Claude-3.5-Sonnet-20241022 $0.2125$ $0.429$ $0.1559$ $0.6889$ $0.2125$ $0.3548$ $0.1559$ $0.1705$ $0.2357$ $0.1705$ $0.5893$ $0.2604$ $0.3118$ $0.5114$ $0.0589$ $0.1705$
Gemini Series Gemini-1.5-Flash $0.2569$ $0.1705$ $0.1559$ $0.1705$ $0.1559$ $0.1705$ $0.1021$ $0.0984$ $0.2041$ $0.1705$ $0.1559$ $0.0984$ $0.1179$ $0.2604$ $0.2125$ $0.1705$
Gemini-2.0-Flash $0.1179$ $0.1705$ $0.3281$ $0.2604$ $0.3584$ $0.2604$ $0.2357$ $0.2952$ $0.4249$ $0.451$ $0.1768$ $0.2952$ $0.2041$ 00 $0.1559$ $0.429$
Gemini-1.5-Pro-002 $0.5103$ $0.6146$ $0.3584$ $0.3937$ $0.4449$ $0.6889$ $0.1021$ $1.135$ $0.4125$ $0.2604$ $0.27$ $0.3548$ $0.2041$ $0.3548$ $0.2041$ $0.2604$
DeepSeek DeepSeek-V3 $0.1021$ $0.6889$ $0.2125$ $0.5208$ $0.4249$ $0.6453$ $0.3536$ $0.2952$ $0.2041$ $0.7811$ $0.8165$ $0.9388$ $1.066$ $0.3548$ $0.3281$ $0.3548$
LLaMA Series LLaMA-3.1-8B-Instruct $0.1559$ $0.743$ $0.1559$ $0.8747$ $0.1021$ $0.2952$ $0.0589$ $0.3548$ $0.2125$ $0.429$ $0.1768$ $0.1968$ $0.2946$ $0.429$ $0.3864$ $0.1705$
LLaMA-3.1-70B-Instruct $0.2946$ $0.3409$ $0.4249$ $0.3548$ $0.3281$ $0.6453$ $0.5137$ $0.2604$ $0.2125$ $0.1705$ $0.1559$ $0.2952$ $0.2041$ $0.0984$ $0.27$ $0.1968$
LLaMA-3.1-405B-Instruct $0.2946$ $0.429$ $0.1559$ $0.1705$ $0.1179$ $0.3937$ $0.1768$ $0.451$ $0.7795$ $0.2604$ $0.3536$ $0.1705$ $0.3536$ $0.1968$ $0.2569$ $0.5208$
LLaMA-3.3-70B-Instruct $0.0589$ $0.3548$ $0.2125$ $0.1968$ $0.2125$ $0.1705$ $0.2125$ 00 $0.2569$ $0.2604$ $0.1179$ $0.1968$ $0.2125$ $0.5208$ $0.0589$ $0.1968$
Qwen Series Qwen2.5-7B-Instruct $0.1559$ $0.0984$ $0.27$ $0.5479$ $0.4677$ $0.1705$ $0.2125$ $0.2604$ $0.1021$ $0.1705$ $0.1559$ $0.0984$ $0.4125$ $0.6818$ $0.2041$ $0.451$
Qwen2.5-14B-Instruct $0.0589$ $0.2952$ $0.2125$ $0.3548$ $0.2041$ $0.5986$ $0.2357$ $0.1968$ $0.3062$ $0.3409$ $0.1021$ $0.1705$ $0.1021$ $0.1968$ $0.1179$ $0.429$
Qwen2.5-32B-Instruct $0.2357$ $0.3548$ $0.3864$ $0.1705$ $0.0589$ $0.0984$ $0.3062$ $0.2952$ $0.1559$ $0.451$ $0.2125$ $0.2604$ $0.0589$ $0.1968$ $0.0589$ $0.2604$
Qwen2.5-72B-Instruct $0.1179$ $0.451$ $0.1559$ $0.0984$ $0.2125$ $0.451$ $0.27$ $0.5479$ $0.4714$ $0.2604$ $0.3536$ $1.028$ $1.327$ $0.8523$ $0.1559$ $0.3937$
Qwen2.5-Coder-32B-Instruct $0.4602$ $0.2604$ $0.4714$ $0.5905$ $0.1768$ $0.858$ $0.5803$ $0.1705$ $0.2041$ $0.7873$ $0.4249$ $0.2952$ 00 $0.3409$ $0.3118$ $0.6889$
Table 9: Standard deviation of LLM performance under four perturbation types for input prediction tasks across Crux in direct inference mode ($N=3$). Red highlights $\sigma>0.5$.
Model Series Model Name VAN ALL MDC MPS MHC
GPT-4 Omni GPT-4o-Mini $0.1021$ $0.8498$ $0.5621$ $0.4602$ $0.2569$
GPT-4o $0.2357$ $0.5303$ $0.27$ $0.1768$ $0.2041$
Claude Series Claude-3.5-Haiku-20241022 $0.2569$ $0.5137$ $0.1021$ $0.2125$ $0.368$
Claude-3.5-Sonnet-20241022 $0.5035$ $0.2569$ $0.1559$ $0.6124$ $0.368$
Gemini Series Gemini-1.5-Flash $0.3062$ $0.1021$ $0.2041$ $0.0589$ $0.0589$
Gemini-2.0-Flash $0.27$ $0.5893$ $0.5035$ $0.5683$ $0.1179$
Gemini-1.5-Pro-002 $0.27$ $0.4082$ $0.4677$ $0.1559$ 00
DeepSeek DeepSeek-V3 $0.3584$ $0.1021$ $0.5803$ $0.6208$ $0.766$
LLaMA Series LLaMA-3.1-8B-Instruct $0.1021$ $0.2125$ $0.4823$ $0.27$ $0.4125$
LLaMA-3.1-70B-Instruct $0.4677$ $0.27$ $0.4249$ $0.27$ $0.4714$
LLaMA-3.1-405B-Instruct $0.2569$ $0.2041$ $0.2569$ $0.5683$ $0.2125$
LLaMA-3.3-70B-Instruct $0.1179$ $0.1021$ $0.27$ $0.2357$ $0.2125$
Qwen Series Qwen2.5-7B-Instruct $0.3118$ $0.3584$ $0.2125$ $0.2125$ $0.6562$
Qwen2.5-14B-Instruct $0.368$ $0.1179$ $0.1768$ $0.5803$ $0.2041$
Qwen2.5-32B-Instruct $0.1021$ $0.3584$ $5890.0$ $0.2125$
Qwen2.5-72B-Instruct $0.6152$ $0.1768$ $0.5137$ $0.0589$ $0.4082$
Qwen2.5-Coder-32B-Instruct $0.27$ $0.4125$ $0.4449$ $0.1559$ $0.2946$
Appendix C Prompts
Our prompt instructions for LLMs follow the format designed by Crux, which is also adopted by LCB, utilizing a two-shot prompt for direct inference and a single-shot prompt for CoT inference to ensure models can follow the instruction and output format effectively. For LRMs, internal reasoning is inherently required before responding. Therefore, applying CoT prompts may lead to redundant reasoning and an unfair comparison with standard LLMs under CoT prompting. Considering the fairness, we use direct inference prompts for RM experiments in §4.3.2.
We provide the complete prompts used for output prediction in both direct and CoT inference (Code 15 and 16) and for input prediction in both direct and CoT inference (Code 17 and 18).
C.1 Output Prediction (Direct)
Code 15: Output Prediction Prompt with Direct Inference.
Given the code snippet:
‘‘‘python
{code}
and the function call with input arguments:
‘‘‘python
{input}
Ensure the provided expression syntax is correct!
Example 1:
Given the code snippet:
‘‘‘python
def f(n):
return n
and the function call with input arguments:
‘‘‘python
f(17)
Example 2:
Given the code snippet:
‘‘‘python
def f(s):
return s + "a"
and the function call with input arguments:
‘‘‘python
f("x9j")
C.2 Output Prediction (CoT)
Code 16: Output Prediction Prompt with CoT-based Inference.
Given the code snippet:
‘‘‘python
{code}
and the function call with input arguments:
‘‘‘python
{input}
Ensure the provided expression syntax is correct!
For example:
Given the code snippet:
‘‘‘python
def f(s):
s = s + s
return "b" + s + "a"
and the input arguments:
‘‘‘python
f("hi")
[THOUGHT]
Lets execute the code step by step:
1. The function f is defined, which takes a single argument s.
5. The return value of the function is therefore "bhihia".
[/THOUGHT]
C.3 Input Prediction (Direct)
Code 17: Input Prediction Prompt with Direct Inference.
Given the code snippet:
‘‘‘python
{code}
and output value:
‘‘‘python
{output}
Example 1:
Given the code snippet:
‘‘‘python
def f(my_list):
count = 0
for i in my_list:
if len(i) % 2 == 0:
count += 1
return count
and output value:
‘‘‘python
3
Example 2:
Given the code snippet:
‘‘‘python
def f(s1, s2):
return s1 + s2
and output value:
‘‘‘python
"banana"
C.4 Input Prediction (CoT)
Code 18: Input Prediction Prompt with CoT-based Inference.
Given the code snippet:
‘‘‘python
{code}
and output value:
‘‘‘python
{output}
For example:
Given the code snippet:
‘‘‘python
def f(x):
return x + 1
and the output value:
‘‘‘python
17
[THOUGHT]
[/THOUGHT]
Appendix D Supplementary Experimental Results
D.1 Impact on Output Prediction
In §4.2, we evaluate the robustness of LLMs on output prediction tasks under direct inference. Table 1 reports the absolute Pass@1 degradation across all perturbations. We use absolute values to highlight the overall impact of each perturbation and conduct cross-perturbation comparisons. For completeness, we also include the relative Pass@1 degradation in Table 10, reflecting the percentage drop from each model’s vanilla performance.
Table 10: Relative Pass@1 degradation of LLMs under all perturbation types for output prediction tasks across Crux and LCB in direct inference ($N=3$). [To Absolute Version]
@@ D.1 Impact on Output Prediction
D.2 Robustness Across Model Scales
In §4.2.3, we compare the robustness against different model scales. To fairly aggregate model robustness across different perturbation types, we compute a weighted average of the relative Pass@1 degradation under four representative perturbations: ALL (structural), MDC and MPS (misleading messages), and MHC (hints). We select ALL as the representative structural perturbation because of the compounding effects of REN, RTF, and GBC captured in §4.2.2, which reflect the most challenging structural setting. Since MDC and MPS stem from the same perturbation strategy and often target overlapping AST node types, we assign them lower weights ($1/8$ each) to avoid over-representing similar effects. Overall, we define a weighted robustness metric with weights of $1/2$, $1/8$, $1/8$, and $1/4$ for ALL, MDC, MPS, and MHC, respectively, and summarize the results in Table 2. For completeness, we also report the results under uniform weighting ($1/4$ for each) in Table 11, which shows similar trends and confirms the stability of our findings in §4.2.3, as well as their absolute degradation versions in Table 12 and 13, respectively.
Table 11: Relative Pass@1 degradation under direct inference, aggregated over Crux and LCB. Weighted average uses: 1/4 ALL, 1/4 MDC, 1/4 MPS, and 1/4 MHC. [To Absolute Version]
@@ D.2 Robustness Across Model Scales
Table 12: Absolute Pass@1 degradation under direct inference, aggregated over Crux and LCB. Weighted average uses: 1/2 ALL, 1/8 MDC, 1/8 MPS, and 1/4 MHC. [To Relative Version]
Model Series Model Name VAN ALL ($1/2$) MDC ($1/8$) MPS ($1/8$) MHC ($1/4$) Average
GPT-4 Omni GPT-4o-Mini 55.77 $-16.34$ $-19.73$ $-13.89$ $-12.98$ $-15.62$
GPT-4o 68.78 $-13.55$ $-13.27$ $-13.55$ $-9.28$ $-12.45$
Claude Series Claude-3.5-Haiku-20241022 57.73 $-16.68$ $-7.14$ $-5.79$ $-21.50$ $-15.33$
Claude-3.5-Sonnet-20241022 72.35 $-16.03$ $-6.28$ $-6.80$ $-18.45$ $-14.26$
Gemini Series Gemini-1.5-Flash 51.73 $-8.57$ $-10.53$ $-16.71$ $-6.25$ $-9.26$
Gemini-2.0-Flash 65.39 $-23.92$ $-23.72$ $-25.91$ $-8.18$ $-20.21$
Gemini-1.5-Pro-002 63.17 $-13.27$ $-16.63$ $-16.05$ $-11.05$ $-13.48$
DeepSeek DeepSeek-V3 67.84 $-14.52$ $-17.54$ $-12.98$ $-12.09$ $-14.10$
LLaMA Series LLaMA-3.1-8B-Instruct 35.47 $-7.90$ $-8.76$ $-7.01$ $-8.84$ $-8.13$
LLaMA-3.1-70B-Instruct 51.89 $-9.77$ $-11.44$ $-15.87$ $-8.37$ $-10.39$
LLaMA-3.1-405B-Instruct 58.72 $-11.02$ $-5.39$ $-8.26$ $-7.61$ $-9.12$
LLaMA-3.3-70B-Instruct 55.64 $-10.14$ $-8.26$ $-9.67$ $-6.36$ $-8.90$
Qwen Series Qwen2.5-7B-Instruct 42.59 $-15.06$ $-21.66$ $-15.32$ $-15.92$ $-16.14$
Qwen2.5-14B-Instruct 48.42 $-17.46$ $-18.09$ $-16.00$ $-6.91$ $-14.72$
Qwen2.5-32B-Instruct 59.89 $-15.22$ $-14.07$ $-14.15$ $-11.81$ $-14.09$
Qwen2.5-72B-Instruct 58.14 $-13.94$ $-8.99$ $-14.44$ $-14.36$ $-13.49$
Qwen2.5-Coder-32B-Instruct 63.10 $-15.27$ $-17.49$ $-13.37$ $-8.99$ $-13.74$
Table 13: Absolute Pass@1 degradation under direct inference, aggregated over Crux and LCB. Weighted average uses: 1/4 ALL, 1/4 MDC, 1/4 MPS, and 1/4 MHC. [To Relative Version]
GPT-4 Omni GPT-4o-Mini 55.77 $-16.34$ $-19.73$ $-13.89$ $-12.98$ $-15.73$
GPT-4o 68.78 $-13.55$ $-13.27$ $-13.55$ $-9.28$ $-12.41$
Claude Series Claude-3.5-Haiku-20241022 57.73 $-16.68$ $-7.14$ $-5.79$ $-21.50$ $-12.78$
Claude-3.5-Sonnet-20241022 72.35 $-16.03$ $-6.28$ $-6.80$ $-18.45$ $-11.89$
Gemini Series Gemini-1.5-Flash 51.73 $-8.57$ $-10.53$ $-16.71$ $-6.25$ $-10.52$
Gemini-2.0-Flash 65.39 $-23.92$ $-23.72$ $-25.91$ $-8.18$ $-20.43$
Gemini-1.5-Pro-002 63.17 $-13.27$ $-16.63$ $-16.05$ $-11.05$ $-14.25$
DeepSeek DeepSeek-V3 67.84 $-14.52$ $-17.54$ $-12.98$ $-12.09$ $-14.28$
LLaMA Series LLaMA-3.1-8B-Instruct 35.47 $-7.90$ $-8.76$ $-7.01$ $-8.84$ $-8.12$
LLaMA-3.1-70B-Instruct 51.89 $-9.77$ $-11.44$ $-15.87$ $-8.37$ $-11.36$
LLaMA-3.1-405B-Instruct 58.72 $-11.02$ $-5.39$ $-8.26$ $-7.61$ $-8.07$
LLaMA-3.3-70B-Instruct 55.64 $-10.14$ $-8.26$ $-9.67$ $-6.36$ $-8.61$
Qwen Series Qwen2.5-7B-Instruct 42.59 $-15.06$ $-21.66$ $-15.32$ $-15.92$ $-16.99$
Qwen2.5-14B-Instruct 48.42 $-17.46$ $-18.09$ $-16.00$ $-6.91$ $-14.61$
Qwen2.5-32B-Instruct 59.89 $-15.22$ $-14.07$ $-14.15$ $-11.81$ $-13.81$
Qwen2.5-72B-Instruct 58.14 $-13.94$ $-8.99$ $-14.44$ $-14.36$ $-12.93$
Qwen2.5-Coder-32B-Instruct 63.10 $-15.27$ $-17.49$ $-13.37$ $-8.99$ $-13.78$
D.3 Effect of CoT Reasoning
In §4.3.1, we analyze the effect of CoT prompting on model robustness using relative degradation aggregated across all perturbations (Table 3). We adopt the relative metric to mitigate the large variance between vanilla performances in direct and CoT inference modes, providing a fair comparison of performance degradation. For completeness, we also report the absolute degradation in Table 14, as well as the relative and absolute degradation on both Crux and LCB in Tables 15 and 16, respectively.
Table 14: Comparison of absolute Pass@1 degradation aggregated over Crux and LCB for output prediction tasks in direct inference ($N=3$) and CoT inference ($N=1$). [To Relative Version]